in a group of 21 people, seven own an ITouch phone. If 9 people are selected at random what is the probability that 5 own an ITouch
Answer by stanbon(48568)

in a group of 21 people, seven own an ITouch phone. If 9 people are selected at random what is the probability that 5 own an ITouch
-----------------------------
# of ways to select 5 owning an ITouch: 7C5 = 7C2 = 7*6/1*2 = 21
---------------------
# of ways to select 4 not owning ITouch: 14C4 = 1001
----------------------
# of ways to select 9 without restriction: 21C9 = 293,930
-----------------------------------------------------------
Probability = [21*1001]/[293,930] = 0.07152
===================================================
Cheers,
Stan H.